# Ctrl-U: Robust Conditional Image Generation via Uncertainty-aware Reward Modeling

### v1

@@ 4.1 Settings
Datasets. Our experiments are conducted using three datasets: ADE20K (Zhou et al., 2017; 2019), COCOStuff (Caesar et al., 2018) and MultiGen-20M dataset (Qin et al., 2024), adhering to the dataset construction principles of Controlnet++ (Li et al., 2024b). More specifically, we utilize the ADE20K and COCOStuff for the segmentation task. For Hed and Lineart conditions, we utilize pretrained models, which is the same as ControlNet++, to generate annotations for the MultiGen-20M dataset. Considering that existing datasets usually include masks for pixels with unknown depth values, we adapt the MultiGen20M depth dataset in a manner similar to the dataset construction method used by ControlNet (Zhang et al., 2023a). For datasets without image captions, we use the captions generated by MiniGPT-4 (Zhu et al., 2023). The training and inference are conducted at a resolution of 512x512 for all datasets and methods.
@@
Evaluation and Metrics. To evaluate semantic segmentation and depth map controls, we adopt mIoU and RMSE as the metrics, respectively. For the Hed edge and Lineart tasks, we calculate SSIM to compare the difference between the extracted control and the ground truth. For Sampling, we employ the UniPC (Zhao et al., 2024b) sampler, implementing 20 denoising steps to generate images using the original text prompts in accordance with ControlNet v1.1 (Zhang et al., 2023a), without incorporating any negative prompts. For comparative methods, we utilized their publicly available code to generate images, ensuring fairness by using the same validation dataset and adhering to their original inference settings.
@@
Implementation Details. In our experiments, we first fine-tune the pre-trained ControlNet model to convergence, using Adam as the optimizer with a learning rate of 1e-5, weight decay of 1e-2, and momentum of 0.9. Then, we use the same optimization settings to perform 10k iterations of uncertainty-aware reward fine-tuning. It is worth noting that we employ an one-step efficient reward strategy (Li et al., 2024b) to enhance training efficiency. Following the settings of previous work, we choose slightly weaker models as the reward model and stronger models for evaluation to guarantee fairness in assessment. We utilize 8 A100 (80G) GPUs for all our experiments. See also the supplementary material for detailed reward model settings.
@@
Baselines. In our evaluation, we mainly compare with competitive methods, including ControlNet++ (Li et al., 2024b), T2I-Adapter (Mou et al., 2024), ControlNet v1.1 (Zhang et al., 2023b), GLIGEN Li et al. (2023), Uni-ControlNet (Zhao et al., 2024a), and UniControl (Qin et al., 2024). These models perform well in controllable text-to-image generation, providing open-source model weights for re-implementation. For fair comparison, all models are tested under the same image conditions and text prompts. Although these methods utilize the user-friendly SD1.5 for controllable text-to-image generation, recent advancements have led to the adoption of SDXL (Podell et al., 2023) by several methods. Accordingly, we also present controllability results for ControlNet-SDXL and T2I-Adapter-SDXL. Since currently there are no officially released models, we deploy the third-party ControlNet-SDXL for the following experiments.
@@ 4.2 Experimental Results
Comparison of Controllability. We present the results on five benchmarks in Table 1. Our uncertainty-aware method Ctrl-U significantly outperforms the previous state-of-the-art method ControlNet++ (Li et al., 2024b) by 6.53% in ADE20K and 8.65% in MultiGen20M depth, respectively. Regarding Hed and Lineart edge, the model with uncertainty regularization has obtained +3.76% and +1.06% increase on SSIM. The most significant improvement is observed in COCO-Stuff for segmentation masks, with an increase of 44.42%. This performance enhancement confirms that reducing the adverse effects of imprecise feedback from the reward model can significantly improve the model controllability. We use the same hyper-parameter settings as those in the ControlNet++ experiments, verifying the effectiveness of our uncertainty-aware reward modeling.
@@
Comparison of CLIP Score. To assess the potential impact on text controllability, we reported the CLIP-Score metrics across various conditional generation tasks in Table 3. We discovered that Ctrl-U generally exhibits comparable or superior CLIP-Score results in most cases, suggesting that our approach not only enhances the controllability of conditional controls but also maintains the original proficiency of text-to-image generation. After communicating with the authors, we re-implemented the CLIP-Score for ControlNet++ (Li et al., 2024b) and marked it with in Table 3.
@@
Human Evaluation. Following ControlNet++ (Li et al., 2024b), we choose four types of conditional controls (segmentation masks, Hed edge, Lineart edge, and depth conditions) for human evaluation. We invite 20 participants to rank each result based on three distinct criteria. According to the average rankings shown in Table 4, users prefer the results produced by our method over those generated by the comparative methods.
@@
Image-Text Alignment 50% 27.5% 22.5% 7.5% 12.5%
@@ 4.3 Ablation Studies and Further Discussion
Impact of Different Timestep Threshold $t_{\text{thre}}$. We investigate the impact of varying timestep threshold $t_{\text{thre}}$ (see Eq. 7) in Table 5(b). We observe that, as the $t_{\text{thre}}$ increases, the scope of our uncertainty-aware reward rectification broadens, leading to improved mIoU. However, while increasing the rectification range can improve alignment between generated images and input conditions, setting the threshold too high will negatively affect image quality, resulting in higher FID scores. We find that the $t_{\text{thre}}=1$ setting strikes an optimal balance, facilitating uncertainty estimation and regularization to adaptively correct the reward learning process.
@@
Table 5: Ablation studies on the ADE20K dataset. We report mIoU, FID and CLIP-score to evaluate controllability and image quality respectively. (a) We show the impact of the interval between $t_{1}$ and $t_{2}$. We find that a short interval limits the diversity of the generated images, making it challenging to estimate uncertainty. In contrast, a long interval impacts accurate uncertainty estimation and thus compromises reward modeling. (b) We study the timestep threshold $t_{\text{thre}}$. We find that the $t_{\text{thre}}=1$ setting strikes an optimal balance, facilitating uncertainty estimation and regularization to adaptively correct the reward learning process. (c) The impact of the regularization weight $\lambda$. With $\lambda=1$, we obtain best results. This setting provides enough model capacity to accurately fit reward feedback while preventing overfitting to inaccurate rewards. (d) The impact of the consistency weight $\mu_{0}$. Considering the balance between diffusion training and uncertainty-aware reward learning, we set $\mu_{0}=0.1$ to ensure enhanced controllability while improving generation quality.
@@
CLIP-score 31.02 30.89 30.84
Figure 3: Qualitative comparisons with different conditional controls on unseen test images.
Figure 4: Ablation study on the uncertainty-aware reward modeling.
@@ 5 Conclusion
Appendix A Appendix
A.1 More Implementation Details.
Dataset Details. We use the ADE20K dataset for segmentation masks, which includes 20,210 images in the training set and 2,000 images in the validation set. The dataset contains 150 types of segmentation annotations, covering most objects found in daily life scenes. Since ADE20K (Zhou et al., 2017; 2019) lacks text prompts, we follow ControlNet++ (Li et al., 2024b) and use MiniGPT-4 (Zhu et al., 2023) to generate image captions by instructing: "Please briefly describe this image in one sentence". Similarly, COCO-Stuff provides segmentation annotations, with 118,287 images in the training set and 5,000 in the validation set. It is challenging to find datasets with accurate annotations for Hed and LineArt edge. To tackle this issue, we adhere to the dataset construction principles of ControlNet (Zhang et al., 2023a) and train our model using the MultiGen20M (Qin et al., 2024) dataset, which includes annotations generated by the model. The images and their corresponding captions are displayed in Fig. 5.
Reward Model Details. We provide details of the models and evaluation metrics in Table 6. Considering that the Hed and Lineart edge extraction models are neural networks without non-differentiable operations, we achieve differentiability by modifying the forward code. We have improved the reward fine-tuning process by the proposed uncertainty-aware reward modeling, resulting in significant enhancement in controllability and generation quality across four control conditions. Our research will expand to encompass additional control scenarios, such as human pose and scribbles.
Table 6: Details of the reward model, evaluation model, and training loss under different conditional controls. ControlNet* indicates that we utilize the same model for condition extraction as employed in ControlNet (Zhang et al., 2023a).
Seg. Mask Depth Edge Hed Edge LineArt Edge
Reward Model (RM) UperNet-R50 DPT-Hybrid ControlNet* ControlNet*
RM Performance ADE20K (mIoU): 42.05 NYU (AbsRel): 8.69 - -
Evaluation Model (EM) Mask2Former DPT-Large ControlNet* ControlNet*
EM Performance ADE20K (mIoU): 56.01 NYU (AbsRel): 8.32 - -
Consistency Loss CrossEntropy Loss MSE Loss MSE Loss MSE Loss
A.2 More Visualization
More visualization results across different conditional controls for our image generation on the validation set are shown in Figures 6789. We observe that our uncertainty-aware framework is capable of generating diverse images that align well with the given inputs across semantic masks, edge, and depth conditions. Attributed to our proposed uncertainty estimation and regularization, the adaptive rectification of inaccurate rewards enhances the diversity of the generated data and its alignment with the given conditions.
Figure 5: Text prompts generated by MiniGPT-4
Figure 6: More visualization results of ours on unseen test images (Segmentation Mask)
Figure 7: More visualization results of ours on unseen test images (Depth Map)
Figure 8: More visualization results of ours on unseen test images (Hed Edge)
Figure 9: More visualization results of ours on unseen test images (LineArt Edge)

### v3

@@ 4.1 Settings
Datasets. Our experiments are conducted using three datasets: ADE20K (Zhou et al., 2017; 2019), COCOStuff (Caesar et al., 2018) and MultiGen-20M dataset (Qin et al., 2023), adhering to the dataset construction principles of Controlnet++ (Li et al., 2024a). More specifically, we utilize the ADE20K and COCOStuff for the segmentation task. For Hed and Lineart conditions, we utilize pretrained models, which is the same as ControlNet++, to generate annotations for the MultiGen-20M dataset. Considering that existing datasets usually include masks for pixels with unknown depth values, we adapt the MultiGen20M depth dataset in a manner similar to the dataset construction method used by ControlNet (Zhang et al., 2023a). For datasets without image captions, we use the captions generated by MiniGPT-4 (Zhu et al., 2023). The training and inference are conducted at a resolution of 512x512 for all datasets and methods.
@@
Evaluation and Metrics. To evaluate semantic segmentation and depth map controls, we adopt mIoU and RMSE as the metrics, respectively. For the Hed edge and Lineart tasks, we calculate SSIM to compare the difference between the extracted control and the ground truth. For Sampling, we employ the UniPC (Zhao et al., 2023b) sampler, implementing 20 denoising steps to generate images using the original text prompts in accordance with ControlNet v1.1 (Zhang et al., 2023a), without incorporating any negative prompts. For comparative methods, we utilized their publicly available code to generate images, ensuring fairness by adhering to their original inference settings.
@@
Implementation Details. In our experiments, we first fine-tune the pre-trained ControlNet model to convergence, using Adam as the optimizer with a learning rate of 1e-5, weight decay of 1e-2, and momentum of 0.9. Then, we use the same optimization settings to perform 10k iterations of uncertainty-aware reward fine-tuning. It is worth noting that we employ an one-step efficient reward strategy (Li et al., 2024a) to enhance training efficiency. Following the settings of previous work, we choose slightly weaker models as the reward model and stronger models for evaluation to guarantee fairness in assessment. See also the supplementary material for detailed reward model settings.
@@
Baselines. In our evaluation, we mainly compare with competitive methods, including ControlNet++ (Li et al., 2024a), T2I-Adapter (Mou et al., 2024), ControlNet v1.1 (Zhang et al., 2023b), GLIGEN Li et al. (2023), Uni-ControlNet (Zhao et al., 2023a), and UniControl (Qin et al., 2023). All of the aforementioned methods have been fine-tuned on datasets across various tasks, and ControlNet++ introduces reward-based post-training on the foundation of ControlNet. These models perform well in controllable text-to-image generation, providing open-source model weights for re-implementation. For fair comparison, all models are tested under the same image conditions and text prompts. Although these methods utilize the user-friendly SD1.5 for controllable text-to-image generation, recent advancements have led to the adoption of SDXL (Podell et al., 2023) by several methods. Accordingly, we also present controllability results for ControlNet-SDXL and T2I-Adapter-SDXL. Since currently there are no officially released models, we deploy the third-party ControlNet-SDXL for the following experiments.
@@ 4.2 Experimental Results
Comparison of Controllability. We present the results on five benchmarks in Table 1. Our uncertainty-aware method Ctrl-U significantly outperforms the previous state-of-the-art method ControlNet++ (Li et al., 2024a) by 6.53% in ADE20K and 8.65% in MultiGen20M depth, respectively. Regarding Hed and Lineart edge, the model with uncertainty regularization has obtained +3.76% and +1.06% increase on SSIM. The most significant improvement is observed in COCO-Stuff for segmentation masks, with an increase of 44.42%. This performance enhancement confirms that reducing the adverse effects of imprecise feedback from the reward model can significantly improve the model controllability. We use the same hyper-parameter settings as those in the ControlNet++ experiments, verifying the effectiveness of our uncertainty-aware reward modeling.
@@
Comparison of CLIP Score. To assess the potential impact on text controllability, we reported the CLIP-Score metrics across various conditional generation tasks in Table 3. We discovered that Ctrl-U generally exhibits comparable or superior CLIP-Score results in most cases, suggesting that our approach not only enhances the controllability of conditional controls but also maintains the original proficiency of text-to-image generation. To ensure a more comprehensive comparison, we re-implemented the CLIP-Score using the official checkpoint for ControlNet++ (Li et al., 2024a), and marked the results with and gray in Table 3.
@@
Human Evaluation. Following ControlNet++ (Li et al., 2024a), we choose four types of conditional controls (segmentation masks, Hed edge, Lineart edge, and depth conditions) for human evaluation.
@@
Image-Text Alignment 50.0% 27.5% 22.5% 7.5% 12.5%
We invite 20 participants to rank each result based on three distinct criteria. According to the average rankings shown in Table 4, users prefer the results produced by our method over those generated by the comparative methods.
@@ 4.3 Ablation Studies and Further Discussion
Impact of Different Timestep Threshold $t_{\text{thre}}$. We investigate the impact of varying timestep threshold $t_{\text{thre}}$ (see Eq. 7) in Table 5(b). We observe that, as the $t_{\text{thre}}$ increases, the scope of our uncertainty-aware reward rectification broadens, leading to improved mIoU. However, while increasing the rectification range can improve alignment between generated images and input conditions, setting the threshold too high will negatively affect image quality, resulting in higher FID scores. We find that the $t_{\text{thre}}=400$ setting strikes an optimal balance, facilitating uncertainty estimation and regularization to adaptively correct the reward learning process.
@@
Table 5: Ablation studies on the ADE20K dataset. We report mIoU, FID and CLIP-score to evaluate controllability and image quality respectively. (a) We show the impact of the interval between $t_{1}$ and $t_{2}$. We find that a short interval limits the diversity of the generated images, making it challenging to estimate uncertainty. In contrast, a long interval impacts accurate uncertainty estimation and thus compromises reward modeling. (b) We study the timestep threshold $t_{\text{thre}}$. We find that the $t_{\text{thre}}=400$ setting strikes an optimal balance, facilitating uncertainty estimation and regularization to adaptively correct the reward learning process. (c) The impact of the regularization weight $\lambda$. With $\lambda=1$, we obtain best results. This setting provides enough model capacity to accurately fit reward feedback while preventing overfitting to inaccurate rewards. (d) The impact of the consistency weight $\mu_{0}$. Considering the balance between diffusion training and uncertainty-aware reward learning, we set $\mu_{0}=0.1$ to ensure enhanced controllability while improving generation quality.
@@
CLIP-score 31.02 31.26 30.84
@@ 5 Conclusion
Figure 3: Qualitative comparisons with different conditional controls on unseen test images. We observe that our generated image preserves condition alignment with good visual quality. Some models do not have open-source weights for Hed or Lineart condition, and thus we skip them.
Figure 4: Ablation study on the uncertainty-aware reward modeling. Here, we show the recovered test image with different denoising timesteps. Since we mitigate the negative impact of noisy rewards, our output maintains consistent semantic conditions with fewer blurred areas.
Acknowledgement
This work is supported by the University of Macau Start-up Research Grant SRG2024-00002-FST and Multi-Year Research Grant MYRG-GRG2024-00077-FST-UMDF.